PART A: Determine the stopping distances for an automobile going a constant initial speed of 95 km/h and human reaction time of 0.30 s , for an acceleration a=−3.6m/s2.

Express your answer using two significant figures.



PART B: Determine the stopping distances for an automobile going a constant initial speed of 95 km/h and human reaction time of 0.30 s , for an acceleration a=−6.5m/s2.

Express your answer using two significant figures.
Answer #1

part A )

initial speed of vehicle , v = 95 km/hr

v = 95 * 1000/3600

v = 26.4 m/s

Now, stopping distance , d = reaction distance + v^2/(2 *a)

d = 26.4 * 0.30 + (26.4)^2/(2 * 3.6)

d = 104.6 m

the stopping distance is 104.6 m

part B)

initial speed of vehicle , v = 95 km/hr

v = 95 * 1000/3600

v = 26.4 m/s

Now, stopping distance , d = reaction distance + v^2/(2 *a)

d = 26.4 * 0.30 + (26.4)^2/(2 * 6.50)

d = 61.5 m

the stopping distance is 61.5 m